Please help me solve this problem: a = 1/2 bh when a = 650 m^2
and b = 20.5 m

Found 2 solutions by checkley77, rfer:
Answer by checkley77(12569)

a = 1/2 bh when a = 650 m^2 and b = 20.5 m
650=20.5h/2
20.5h=650*2
20.5h=1,300
h=1,300/20.5
h=63.414634 is the height.

Answer by rfer(10417)

A=.5bh
650=.5(20.5)h
650=10.25h
divide both sides by 10.25
h=63.41m